(1 point)Several Values of a continuous function f(x) are given below: (1 point) Several values of a continuous function f(x) are given below: 2= f(x) = 1 9 2 -12 3 5 4 5 Find (Si (f (t)) dt)|g_1
Answer #1

x = 1 2 3 f(x) = 9.-125 4 a 5 Now, fl stuu it + f() m (M) - f (1) au f - otfrog - 0 = fun sol the firme) (x=1 - f (1)